# Peculiar dark matter halos inferred from gravitational lensing as a manifestation of modified gravity

### v2

@@
If modified gravity holds, but the weak lensing analysis is done in the standard way, one finds that dark matter halos have peculiar shapes, not following the standard Navarro-Frenk-White profiles, and are fully predictable from the distribution of baryons. Here we study in detail the distribution of the apparent dark matter around point masses, which approximate galaxies and galaxy clusters, and their pairs for the QUMOND MOND gravity, taking an external gravitational acceleration $g_{e}$ into account. At large radii, the apparent halo of a point mass $M$ is shifted against the direction of the external field. When averaged over all lines-of-sight, the halo has a hollow center, and denoting the by $a_{0}$ the MOND acceleration constant, its density behaves like $\rho(r)=\sqrt{Ma_{0}/G}/(4\pi r^{2})$ between the galacticentric radii $\sqrt{GM/a_{0}}$ and $\sqrt{GMa_{0}}/g_{e}$, and like $\rho\propto r^{-7}G^{2}M^{3}a_{0}^{3}/g_{e}^{5}$ further away. Between a pair of point masses, there is a region of a negative apparent dark matter density, whose mass can exceed the baryonic mass of the system. The density of the combined dark matter halo is not a sum of the densities of the halos of the individual points. The density has a singularity near the zero-acceleration point, but remains finite in projection. We compute maps of the surface density and the lensing shear for several configurations of the problem, and derive formulas to scale them to further configurations. In general, for a large subset of MOND theories in their weak field regime, for any configuration of the baryonic mass $M$ with the characteristic size of $d$, the total lensing density scales as $\rho({\@vec{x}})=\sqrt{Ma_{0}/G}d^{-2}f\left(\@vec{\alpha},\@vec{x}/d,g_{e}d/\sqrt{GMa_{0}}\right)$, where the vector ${\alpha}$ describes the geometry of the system. Detecting the difference between QUMOND and cold dark matter halos appears to be possible with the existing instruments.
@@ 1 Introduction
The missing mass problem has not been solved decisively yet. The usual solution is to assume the validity of Newtonian dynamics and general relativity and to postulate the existence of dark matter (see, e.g., Salucci, 2019 for a review). While many types of dark matter have been proposed, here we for concreteness focus on the cold dark matter and the $\Lambda$CDM cosmological model. In the $\Lambda$CDM model, many well-known observations, particularly on the cosmological scales, can be explained well (cosmic microwave background, baryon acoustic oscillations, distance-luminosity relation for type II supernovae, etc.). However, many questions remain unexplained, such as the mismatch between the Hubble constant in the early and late Universe (Verde et al., 2019), large-scale cosmic flows (Peery et al., 2018), emptiness of the Local Void (Peebles & Nusser, 2010; Haslbauer et al., 2020), existence of the planes of satellites (Pawlowski et al., 2019), among others. The internal properties of galaxies cannot be predicted well from the first principles and are even hard to reproduce in simulations (Brooks & Christensen, 2016; Oman et al., 2015; Bullock & Boylan-Kolchin, 2017), while simple, tight scaling relations between the baryonic and dark matter content are observed (Famaey & McGaugh, 2012; McGaugh et al., 2016).
@@
In modified gravity theories which do not involve dark matter, we introduce the concept of “phantom dark matter” (PDM, Milgrom, 1986). It is a mathematical construct that denotes the dark matter that would lead in the Newtonian gravity to the observed gravitational field while actually a modified gravity theory holds true. The distribution of PDM can be very unusual or even impossible from the point of view of the Newtonian gravity. For example, in MOND the PDM around a point mass has a zero density in the vicinity of the point, but at larger radii it mimics an isothermal halo. This is in stark contrast to the Navarro-Frenk-White dark matter halo expected by the theory of $\Lambda$CDM, which has an infinite density in its center (Navarro et al., 1996). For some configurations of baryonic matter, the density of PDM can even be negative in some regions of space. Milgrom (1986) identified a few specific mass configurations that give rise to negative density of PDM: a pair of point masses, a disk with a hole in its center, an inhomogeneous thin galactic disk, and an infinitesimal mass embedded in a homogeneous external gravitational field. Oria et al. (2021) made maps of the regions of the negative PDM for the Local Volume. The distribution of PDM around an object depends, in general, on the external field imposed on it by other objects. Also, it can strongly depend on the particular MOND theory (Milgrom, 2023).
@@
In the current paper, we investigate in detail PDM halos of point masses and their pairs and the influence of the EFE on them. We also present the predicted maps of gravitational lensing around them. In particular, we do this for the QUMOND modified gravity version of MOND (Milgrom, 2010). The point masses are assumed to represent galaxies, galaxy groups and clusters as a zeroth order approximation. We investigate whether the special effects of MOND distinguishing it from the $\Lambda$CDM cosmology are within the reach of the current instruments. The topic is particularly timely since the recently launched Euclid satellite is expected to revolutionize the field of weak lensing. It will survey about 1/3 of the sky with an angular resolution comparable to the Hubble Space Telescope. The multitude of the photometric filters will not only allow determining the redshifts of the lensed galaxies, but also of the accurate stellar masses of the lenses, that are needed for predicting the gravitational field in MOND. Unlike the distinguishing test of MOND and particle dark matter based on dynamical measurements, weak lensing does not impose any assumptions on the dynamics of the investigated objects (such as being in virial equilibrium) and is insensitive to many of the uncertain aspects of galaxy evolution.
The paper is organized as follows. We explain how to calculate the PDM density in QUMOND in Sect. 2. Then we remind the weak lensing formalism in general in Sect. 3. We progress to the case of a single point mass in Sect. 5. We discuss the case when there is no external field (Sect. 5.1), external field with a known direction (Sect. 5.2) and external field with an unknown direction (Sect. 5.3). We also compare the QUMOND predicted PDM halo to a dark halo in the $\Lambda$CDM cosmology in Sect. 5.4. Section 6 is devoted to a pair of point masses, observed either perpendicularly to the connecting line (Sect. 6.2), or along it (Sect. 6.3). The external field is treated in an approximate way and, therefore, this approximation is justified in detail in Sect. 6.4. In Sect. 7, we estimate whether the deviation of QUMOND from Newtonian gravity can be achieved with the current instruments and we find that it indeed seems feasible. We summarize our findings in Sect. 8.
@@ 2 Calculating the effective distribution of the lensing mass in modified gravity
In Appendix A we derive general functional forms for the density and surface density of the PDM in the deep-MOND regime of any MOND theory in which the concept of PDM makes sense and does not introduce any other new constants of physics than $a_{0}$. Namely, if $M$ denotes the total baryonic mass of the system, $d$ its characteristic size, $g_{e}$ the magnitude of the external field, and ${\alpha}$ a vector of dimensionless parameters that describe the geometry of the problem, then:
@@ 3 Methods to predict the maps of gravitational lensing
[…]
@@
where $\delta_{ij}$ is Kronecker’s symbol. If we calculate the eigenvectors ${e_{+}}$ and ${e_{-}}$ and the corresponding eigenvalues $\lambda_{+}$ and $\lambda_{-}$ of the inverse of the Jacobian matrix, $A^{-1}$, we learn to what ellipse a circular source is deformed into. The major axis of the ellipse points in the direction of ${e_{+}}$ that corresponds to the larger eigenvalue $\lambda_{+}$. The ratio of the eigenvalues is equal to the ratio of the axes of the ellipse. The complex ellipticity (one of the two common definitions which is related more directly to observations, see Bartelmann & Schneider, 2001) is defined as
@@ 5.1 Zero external field
Let us review the case of point mass that is in a zero external field. Gravitational acceleration ${g}$ of any spherically symmetric object in QUMOND can be calculated as (Milgrom, 2010):
${g}=\@vec{g_{N}}\nu\left(g_{N}/a_{0}\right).$ (19)
By applying the Poisson equation to this field, one gets the density of the PDM. From Eq. 19, one can see that ${g}\approx\@vec{g_{N}}$ in the region where $g_{N}\gg a_{0}$, that is well below the so-called MOND transitional radius
@@
Milgrom & Sanders (2008) explored the surface distribution of PDM for a variety of circularly symmetric sources and interpolation functions. They found that MOND predicts rings in the surface density of PDM, unless the interpolation function reaches the high-acceleration limit too slowly. The maximum surface density of the ring is on the order of the constant $\Sigma_{0}=a_{0}/G=8.6\times 10^{14}\,M_{\sun}\,$pc-2 and it is reached approximately at the distance of $r_{\mathrm{M}}$.
@@ 5.3 Unknown direction of the external field
    The profile of the density or surface density of the PDM halo is strongly affected by the external field only in the regions that are further from the point mass than $r_{\mathrm{ef}}$. We need to measure surface density with a precision better than about $10^{12}\,M_{\sun}\,$pc-2. The differences in the $r_{\mathrm{ef}}$ radii are in the order of hundreds of kpc for the range of external field magnitudes considered here.
@@ 5.4 Comparison to an NFW halo
Between the radii $r_{\mathrm{M}}$ and $r_{\mathrm{ef}}$, the density of PDM is roughly like of an isothermal sphere, given by Eq. 21. This is in contrast with the predictions of the $\Lambda$CDM cosmology, where dark matter halos are expected to have a density profile similar to the Navarro-Frenk-White (NFW) profile (Navarro et al., 1996). We thus could discriminate between the $\Lambda$CDM model and MOND if we could decide what profile of (phantom) dark halos galaxies have. We quantified the difference for one representative galaxy. We chose the baryonic mass of the galaxy as $M=4\times 10^{10}\,M_{\sun}$, similar to the characteristic mass $M_{*}$ of the galaxy mass function (Kawinwanichakij et al., 2020). Then we assigned a dark matter halo to the galaxy making use of the mean stellar-to-halo relation (Behroozi et al., 2013) and halo mass-concentration relation (Diemer & Kravtsov, 2015) at zero redshift. This yielded the virial mass of the halo of $1.6\times 10^{12}\,M_{\sun}$ and a virial radius of 310 kpc. We studied the radial range only up to the virial radius (see, e.g., Diemer & Kravtsov, 2015 for a more detailed behaviour of $\Lambda$CDM halos at large radii).
@@
The figure shows that for detecting the deviation of the PDM halo from the NFW halo need to detect a difference in surface densities on the order of $10^{13}\,M_{\sun}\,$pc-2 or in tangential shears on the order of $10^{-3}$ in the radial range of tens of arcsec. We will discuss whether such a measurement is feasible in Sect. 7. The deviation of the PDM halo from the NFW halo beyond about 100 kpc depends strongly on the magnitude of the external field.
@@ 6.1 Distribution of PDM around two point masses
In this section, we will investigate the distribution of PDM around a pair of point masses in a zero external field. Let us denote the masses of the two points by $M_{1}$ and $M_{2}$, such that $M_{1}\geq M_{2}$, and their distance by $d$. We will work in the cylindrical coordinate system that has its axial coordinate, $z$ coincident with the connecting line of the point masses. We denote the radial coordinate by $r$. Let the point mass $M_{1}$ be located at $z=0$ and $M_{2}$ at $z=-d$. The density $\rho_{\mathrm{ph}}$ or projected surface density $\Sigma_{\mathrm{ph}}$ of the PDM at the position ${x}$ can depend only on $G,a_{0},M_{1},M_{2},d$ and ${x}$. For dimensional reasons, $\rho_{\mathrm{ph}}$ and $\Sigma_{\mathrm{ph}}$ must be given by expressions of the forms:
@@
$M_{\mathrm{tot}}=M_{1}+M_{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \eta=\frac{M_{2}}{M_{1}},\leavevmode\nobreak\ \leavevmode\nobreak\ \textrm{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \theta=\frac{\sqrt{GM_{\mathrm{tot}}/a_{0}}}{d}.$ (30)
@@
$-4\pi G\rho_{\mathrm{ph}}=\partial_{i}\left[\nu g_{N}^{i}\right]=\partial_{i}(\nu)g_{N}^{i}+\nu\partial_{i}g_{N}^{i}=\frac{\nu^{\prime}}{a_{0}}\partial_{i}(g_{N}^{j}g_{N,j})^{1/2}g_{N}^{i}.$ (31)
@@
We characterize the shape of the negative PDM region by its maximum radial extent $r_{\mathrm{max}}$ and its minimum and maximum $z$ coordinate with respect to the position of the mass $M_{2}$ (i.e., $z_{\mathrm{min}}=0$ indicates that the minimum $z$ coordinate is reached at $z=-d$). It follows from the previous paragraph that any linear dimension of this region, $x$, depends only on $G,M_{1},M_{2}$, and $d$. A dimensional analysis then indicates that $x/d=f\left(\eta\right)$. We derived the function numerically for $r_{\mathrm{max}}$, $z_{\mathrm{min}}$ and $z_{\mathrm{max}}$ and is shown in Fig. 6. The value of $z_{\mathrm{min}}$ becomes negative for $\eta<3.91\times 10^{-3}$.
Next, let us focus on the mass and characteristic surface density of the region of negative PDM. A dimensional analysis gives that the mass of the region, $M_{\mathrm{neg}}$, satisfies $M_{\mathrm{neg}}/M_{\mathrm{tot}}=f(\eta,\theta)$. We calculated the function numerically in the top panel of Fig. 7. It is interesting that the mass of the negative region can be comparable, or even higher than the total baryonic mass of the system. This happens preferably for systems consisting of two similar masses in deep-MOND regime. For example, for two galaxy clusters with a mass of $10^{13}\,M_{\sun}$ located 151515\,15Mpc apart, the $\theta$ parameter is about $10^{-2}$ and therefore the region of negative PDM would have a mass of a galaxy cluster.
@@
$\begin{split}&\rho_{\mathrm{ph}}\sim\frac{\left(\!\!\sqrt{M_{1}}+\sqrt{M_{2}}\right)^{2}}{8\pi d^{3/2}(M_{1}M_{2})^{1/4}}\sqrt{\frac{a_{0}}{G}}\frac{8(z-z_{0})^{2}-r^{2}}{\left[4(z-z_{0})^{2}+r^{2}\right]^{5/4}}\\ &\textrm{for}\leavevmode\nobreak\ \leavevmode\nobreak\ (r,z)\rightarrow(0,z_{0}).\end{split}$ (32)
@@ 6.2.3 Results
The magnitude of the dip is generally a factor of a few. For the fiducial model, the dip occurs at the surface density level of about $10^{13}\,M_{\sun}\,$pc-2. This surface density grows the strongest with the $\theta$ parameter. This is because the density of PDM is the highest near the MOND transitional radius [^1] The size of the MOND transitional radius is negligible in the images. (Sect. 5) and the points become closer to each other with decreasing $\theta$. We note that even if the PDM density diverges near the zero acceleration point (Eq. 32), the surface density does not, as proved analytically in Appendix B.
@@
On the other hand, the superposition surface density is less than the true PDM surface density in the outer parts of the halo. This is caused by the external field in which the point masses are embedded both: the combined PDM halo of the pair of point masses is truncated by the EFE at a larger radius, than the PDM halos of the single point masses that form the superposition halo. We note that in the external parts of the halo, the difference is relatively constant, having the value of around $10^{12}\,M_{\sun}$pc-2.
In the theories postulating particle dark matter and Newtonian gravity, the real density and superposition density will differ too, but probably in a different way. During the first approach of the two halos toward each other, the tidal force will bulge the halos along their connection line, producing a surplus of dark matter in between of the galaxies. Later, once the halos experience tidal stripping and compactification, the result is more difficult to predict. Nevertheless, (Pawlowski et al., 2017) inspected the distribution of satellites around isolated pairs of galaxies in a $\Lambda$CDM simulation and found that they tend to be lopsided toward each other. In contrast, Fig. 14 shows that the PDM halos in QUMOND appear “squashed” in the region between the point masses. Here, the density of PDM is reduced compared to the superposition density, see Fig. 16. At least this is true for the surface densities. Figure 5 shows that the volume density contours of PDM halos have more intricate shapes.

### v3

@@
If modified gravity holds, but the weak lensing analysis is done in the standard way, one finds that dark matter halos have peculiar shapes, not following the standard Navarro-Frenk-White profiles, and are fully predictable from the distribution of baryons. Here we study in detail the distribution of the apparent dark matter around point masses, which approximate galaxies and galaxy clusters, and their pairs for the QUMOND MOND gravity, taking an external gravitational acceleration $g_{e}$ into account. At large radii, the apparent halo of a point mass $M$ is shifted against the direction of the external field. When averaged over all lines-of-sight, the halo has a hollow center, and denoting the by $a_{0}$ the MOND acceleration constant, its density behaves like $\rho(r)=\sqrt{Ma_{0}/G}/(4\pi r^{2})$ between the galacticentric radii $\sqrt{GM/a_{0}}$ and $\sqrt{GMa_{0}}/g_{e}$, and like $\rho\propto r^{-7}G^{2}M^{3}a_{0}^{3}/g_{e}^{5}$ further away. Between a pair of point masses, there is a region of a negative apparent dark matter density, whose mass can exceed the baryonic mass of the system. The density of the combined dark matter halo is not a sum of the densities of the halos of the individual points. The density has a singularity near the zero-acceleration point, but remains finite in projection. We compute maps of the surface density and the lensing shear for several configurations of the problem, and derive formulas to scale them to further configurations. In general, for a large subset of MOND theories in their weak field regime, for any configuration of the baryonic mass $M$ with the characteristic size of $d$, the total lensing density scales as $\rho({\@vec{x}})=\sqrt{Ma_{0}/G}d^{-2}f\left(\@vec{\alpha},\@vec{x}/d,g_{e}d/\sqrt{GMa_{0}}\right)$, where the vector $\@vec{\alpha}$ describes the geometry of the system. Detecting the difference between QUMOND and cold dark matter halos appears to be possible with the existing instruments.
@@ 1 Introduction
[…]
@@
[…]
@@
[…]
@@ 2 Calculating the effective distribution of the lensing mass in modified gravity
In Appendix A we derive general functional forms for the density and surface density of the PDM in the deep-MOND regime of any MOND theory in which the concept of PDM makes sense and does not introduce any other new constants of physics than $a_{0}$. Namely, if $M$ denotes the total baryonic mass of the system, $d$ its characteristic size, $g_{e}$ the magnitude of the external field, and $\@vec{\alpha}$ a vector of dimensionless parameters that describe the geometry of the problem, then:
@@ 3 Methods to predict the maps of gravitational lensing
In these formulas ${\@vec{\theta}}=(\theta_{1},\theta_{2})$ is a vector, expressed in the units of radians, in a Cartesian system that touches the celestial sphere at the position of the lens from the point of view of the observer. For circularly symmetric lenses we match the origin of the coordinate system with the center of the symmetry. The symbol $\Delta$ stands for the two-dimensional Laplace operator with respect to the coordinates $\@vec{\theta}$. The surface density $\Sigma$ in Eq. 6 is given by integration of the volume density $\rho$ of the lens along the line of sight $z$. In the context of this paper, $\rho$ is the sum of the density of the baryonic matter and the PDM. The constant $\Sigma_{\mathrm{cr}}$ is called the critical surface density. In Eq. 7, $D_{L}$ denotes the angular-diameter distance of the lens from the observer, $D_{S}$ the angular-diameter distance of the source from the observer, and $D_{LS}$ the angular-diameter distance of the source for an observer at the lens at the cosmic time when the photon passes the lens. For a point mass $M$ (without a PDM halo) we have
@@
where $\delta_{ij}$ is Kronecker’s symbol. If we calculate the eigenvectors $\@vec{e_{+}}$ and $\@vec{e_{-}}$ and the corresponding eigenvalues $\lambda_{+}$ and $\lambda_{-}$ of the inverse of the Jacobian matrix, $A^{-1}$, we learn to what ellipse a circular source is deformed into. The major axis of the ellipse points in the direction of $\@vec{e_{+}}$ that corresponds to the larger eigenvalue $\lambda_{+}$. The ratio of the eigenvalues is equal to the ratio of the axes of the ellipse. The complex ellipticity (one of the two common definitions which is related more directly to observations, see Bartelmann & Schneider, 2001) is defined as
@@ 5.1 Zero external field
Let us review the case of point mass that is in a zero external field. Gravitational acceleration g\@vec{g} of any spherically symmetric object in QUMOND can be calculated as (Milgrom, 2010):
$\@vec{g}=\@vec{g_{N}}\nu\left(g_{N}/a_{0}\right).$ (19)
By applying the Poisson equation to this field, one gets the density of the PDM. From Eq. 19, one can see that $\@vec{g}\approx\@vec{g_{N}}$ in the region where $g_{N}\gg a_{0}$, that is well below the so-called MOND transitional radius
@@
Milgrom & Sanders (2008) explored the surface distribution of PDM for a variety of circularly symmetric sources and interpolation functions. They found that MOND predicts rings in the surface density of PDM, unless the interpolation function reaches the high-acceleration limit too slowly. The maximum surface density of the ring is on the order of the constant $\Sigma_{0}=a_{0}/G=8.6\times 10^{14}\,M_{\sun}\,$Mpc-2 and it is reached approximately at the distance of $r_{\mathrm{M}}$.
@@ 5.3 Unknown direction of the external field
    The profile of the density or surface density of the PDM halo is strongly affected by the external field only in the regions that are further from the point mass than $r_{\mathrm{ef}}$. We need to measure surface density with a precision better than about $10^{12}\,M_{\sun}\,$Mpc-2. The differences in the $r_{\mathrm{ef}}$ radii are in the order of hundreds of kpc for the range of external field magnitudes considered here.
@@ 5.4 Comparison to an NFW halo
[…]
@@
The figure shows that for detecting the deviation of the PDM halo from the NFW halo need to detect a difference in surface densities on the order of $10^{13}\,M_{\sun}\,$Mpc-2 or in tangential shears on the order of $10^{-3}$ in the radial range of tens of arcsec. We will discuss whether such a measurement is feasible in Sect. 7. The deviation of the PDM halo from the NFW halo beyond about 100 kpc depends strongly on the magnitude of the external field.
@@ 6.1 Distribution of PDM around two point masses
In this section, we will investigate the distribution of PDM around a pair of point masses in a zero external field. Let us denote the masses of the two points by $M_{1}$ and $M_{2}$, such that $M_{1}\geq M_{2}$, and their distance by $d$. We will work in the cylindrical coordinate system that has its axial coordinate, $z$ coincident with the connecting line of the point masses. We denote the radial coordinate by $r$. Let the point mass $M_{1}$ be located at $z=0$ and $M_{2}$ at z=dz=-d. The density $\rho_{\mathrm{ph}}$ or projected surface density $\Sigma_{\mathrm{ph}}$ of the PDM at the position x\@vec{x} can depend only on $G,a_{0},M_{1},M_{2},d$ and x\@vec{x}. For dimensional reasons, $\rho_{\mathrm{ph}}$ and $\Sigma_{\mathrm{ph}}$ must be given by expressions of the forms:
@@
$M_{\mathrm{tot}}=M_{1}+M_{2},\penalty 10000\ \penalty 10000\ \penalty 10000\ \eta=\frac{M_{2}}{M_{1}},\penalty 10000\ \penalty 10000\ \textrm{and}\penalty 10000\ \penalty 10000\ \theta=\frac{\sqrt{GM_{\mathrm{tot}}/a_{0}}}{d}.$ (30)
@@
$4\pi G\rho_{\mathrm{ph}}=\partial_{i}\left[\nu g_{N}^{i}\right]=\partial_{i}(\nu)g_{N}^{i}+\nu\partial_{i}g_{N}^{i}=\frac{\nu^{\prime}}{a_{0}}\partial_{i}(g_{N}^{j}g_{N,j})^{1/2}g_{N}^{i}.$ (31)
@@
We characterize the shape of the negative PDM region by its maximum radial extent $r_{\mathrm{max}}$ and its minimum and maximum $z$ coordinate with respect to the position of the mass $M_{2}$ (i.e., $z_{\mathrm{min}}=0$ indicates that the minimum $z$ coordinate is reached at z=dz=-d). It follows from the previous paragraph that any linear dimension of this region, $x$, depends only on $G,M_{1},M_{2}$, and $d$. A dimensional analysis then indicates that $x/d=f\left(\eta\right)$. We derived the function numerically for $r_{\mathrm{max}}$, $z_{\mathrm{min}}$ and $z_{\mathrm{max}}$ and is shown in Fig. 6. The value of $z_{\mathrm{min}}$ becomes negative for $\eta<3.91\times 10^{-3}$.
Next, let us focus on the mass and characteristic surface density of the region of negative PDM. A dimensional analysis gives that the mass of the region, $M_{\mathrm{neg}}$, satisfies $M_{\mathrm{neg}}/M_{\mathrm{tot}}=f(\eta,\theta)$. We calculated the function numerically in the top panel of Fig. 7. It is interesting that the mass of the negative region can be comparable, or even higher than the total baryonic mass of the system. This happens preferably for systems consisting of two similar masses in deep-MOND regime. For example, for two galaxy clusters with a mass of $10^{13}\,M_{\sun}$ located 1515\,Mpc apart, the $\theta$ parameter is about $10^{-2}$ and therefore the region of negative PDM would have a mass of a galaxy cluster.
@@
$\begin{split}&\rho_{\mathrm{ph}}\sim\frac{\left(\!\!\sqrt{M_{1}}+\sqrt{M_{2}}\right)^{2}}{8\pi d^{3/2}(M_{1}M_{2})^{1/4}}\sqrt{\frac{a_{0}}{G}}\frac{8(z-z_{0})^{2}-r^{2}}{\left[4(z-z_{0})^{2}+r^{2}\right]^{5/4}}\\ &\textrm{for}\penalty 10000\ \penalty 10000\ (r,z)\rightarrow(0,z_{0}).\end{split}$ (32)
@@ 6.2.3 Results
The magnitude of the dip is generally a factor of a few. For the fiducial model, the dip occurs at the surface density level of about $10^{13}\,M_{\sun}\,$Mpc-2. This surface density grows the strongest with the $\theta$ parameter. This is because the density of PDM is the highest near the MOND transitional radius [^1] The size of the MOND transitional radius is negligible in the images. (Sect. 5) and the points become closer to each other with decreasing $\theta$. We note that even if the PDM density diverges near the zero acceleration point (Eq. 32), the surface density does not, as proved analytically in Appendix B.
@@
On the other hand, the superposition surface density is less than the true PDM surface density in the outer parts of the halo. This is caused by the external field in which the point masses are embedded both: the combined PDM halo of the pair of point masses is truncated by the EFE at a larger radius, than the PDM halos of the single point masses that form the superposition halo. We note that in the external parts of the halo, the difference is relatively constant, having the value of around $10^{12}\,M_{\sun}$Mpc-2.
[…]